# Search for high energy 5.5 MeV solar axions with the complete Borexino dataset

### v1

@@ 4.2 Response function to axion reactions
The signature of these four reactions discussed above is a $5.5$ MeV peak. The Monte Carlo (MC) method based on the GEANT4 framework was used to simulate the Borexino response to electrons and $\gamma$-quanta produced by axion interactions Ago18 . The simulations have taken into account the effect of ionization quenching, Cerenkov radiation and small non-linearity induced by the energy dependence on the event position.
@@
Table 1: The fraction $\epsilon$ of axion events surviving data selection procedure. CC - Compton axion to photon conversion, $A+e\rightarrow e+\gamma$; AE - axioelectric effect, $A+e+Z\rightarrow e+Z$; PC - Primakoff conversion on nuclei, $A+{{}^{12}\rm{C}}\rightarrow\gamma+{{}^{12}\rm{C}}$.
@@
Figure 3: Simulated responses to axion interactions in the Borexino IV: 1 – axioelectric effect ($5.49$ MeV electrons), 2 – Compton axion to photon conversion (electrons and $\gamma$-quanta), 3 – decay $A\rightarrow 2\gamma$, 4 – Primakoff conversion ($5.49$ MeV $\gamma$-quanta). The response functions are normalized to single axion interaction or decay in $5$ m radius sphere.
@@ 4.3 Data selection
The Borexino detector carried out measurements from May 15, 2007 to October 7, 2021. The aim of data selection is to provide maximum exposure for the desired study with minimum background contribution. Background composition of the Borexino experiment was carefully studied in the course of many years of research. In the current analysis we aim to the energy region around of 5.5 MeV. The background components that have to be taken into account are the short-lived cosmogenic beta-sources (${}^{12}\rm{B}$, 8He, 9C, 9Li etc.), decays of internal 208Tl, external high-energy gamma produced due to neutron radiative capture in the SSS and solar neutrinos produced in the 8B decay.
@@
This system of cosmogenic veto allows for maximal suppression of the cosmogenic background for the price of $22.8$% exposure loss.
Whilst the internal 208Tl and the solar neutrinos are uniform in the bulk of the detector and can not be suppressed or discriminated, the external high-energy gammas can be discriminated. Since these events count rate decreases exponentially towards the center of the detector, they can be suppressed by setting up a fiducial volume. In this analysis the fiducial volume was defined by minimal distance to the inner vessel of $0.75$ m and had the mass of around $145$ tons. Moreover, these events were simulated with a full MC simulation following the detector evolution and this simulation was used for a so-called multivariate fit that takes into account the radial distribution providing a discrimination between external and internal spectrum components.
Figure 4: Energy spectra of the events and effect of the selection cuts. From top to bottom: (1) raw spectrum; (2) with $2$ ms muon veto cut; (3) final spectrum after application of the cosmogenic veto and FV cut. The bumps at $0.4$, $1.3$, $2.4$ and $3.3$ MeV are caused by decays of 210Po, 11C, external $\gamma$’s from 208Tl and internal 208Tl, respectively.
The experimental energy spectrum from Borexino in the range up to 6600 p.e. ($\simeq 12$ MeV), containing 579.3 kton days of data, is shown in Fig.4. At energies below 3 MeV, the spectrum is dominated by 2.6 MeV $\gamma$’s from the $\beta$-decay of 208Tl in the PMTs and in the SSS. The bumps in the spectrum observed at $200$ p.e., $700$ p.e., $1300$ p.e. and $1800$ p.e. are caused by $\alpha$-decays of 210Po, $\beta^{+}$-decays of $\rm{{}^{11}C}$, external $\gamma^{\prime}s$ from $\rm{{}^{208}Tl}$ and internal $\beta^{-}$-decays of $\rm{{}^{208}Tl}$, respectively (more details e.g. in Ago19 ).
@@ 5.1 Analysis of the spectrum
Figure 5 shows the observed the Borexino energy spectrum in the ($4.4-6.8$) MeV ($(2200-3400)$ p.e.) range in which the axion peaks might appear. The spectrum $N(E)$ was fitted by the sum of four main components: the $\rm{{}^{208}Tl}$ decays inside the scintillator $N_{int}$, the simulated spectrum of external backgrounds caused by consequent $(\alpha,n)$ reactions followed by neutron captures $(n,\gamma)$ on Ni, Fe, Cr nuclei contained in the SSS constructions $N_{ext}$, the solar $\rm{{}^{8}B}$-neutrino-electron scattering events $N_{8B\nu}$, and the simulated spectrum of axion response $N_{A}$.
@@
Figure 5: Fit of the Borexino energy spectrum in the ($4.4-6.8$) MeV range. Curve 1 is the detector response function for axio-electric effect on carbon atoms and corresponds to $90$% c.l. upper limit ($S=5.6$ events). The fit is complement to the radial fit on Fig. 6.
Figure 6: The fitted Borexino radial spectrum. The curve 1 corresponds to the uniform part of the spectrum whilst the curve 2 is the radial distribution of the simulated external background events. The decline in the distribution of uniform events is caused by the fiducial volume non-sphericity. The fit is complementary to the radial fit on Figure 5.
The limits obtained ($S^{\rm lim}_{avg}\simeq 1.7\times 10^{-3}$ counts/(100 t day) at 90% c.l.) are very low, e.g. $\sim 8\times 10^{4}$ times lower than expected number of events from solar $pp-$ neutrino (135 counts/(100 t day)). The upper limits on the number of events with energy 5.5 MeV constrain the product of axion flux $\Phi_{A}$ and the interaction cross section with electron, proton or carbon nucleus $\sigma_{A-e,p,C}$ via
@@
where $N_{e,p,C}$ is the number of electrons, protons and carbon nuclei and $T$ is the measurement time, the product of these factors corresponds to the statistics of 579.3 kton×$\times$×day. The individual rate limits are:
@@
5.2 Limits on $g_{A\gamma}\times g_{3AN}$-coupling
@@
The relationship between the isovector constant $g_{3AN}$ and the axion mass $m_{A}$ allows to constraint the $g_{A\gamma}$ depending on $m_{A}$. The three cases mentioned above were considered - the KSVZ-model with $E/N=0$ and the DFSZ-model for the angles $\tan\beta=140$ and $\tan\beta=0.25$ which are the boundary values of the allowed unknown angle $\beta$.
The excluded regions are inside three contours 1 in Fig. 7 (90 % c.l.). The curves from bottom to top correspond to DFSZ ($\tan\beta=140$), KSVZ and DFSZ ($\tan\beta=0.25$), respectively. For higher values of $g_{A\gamma}^{2}m_{A}^{3}$ axions decay before they reach the detector, while for lower $g_{A\gamma}^{2}m_{A}^{3}$ the probability of axion decay inside the Borexino volume is too low. The limits on $g_{A\gamma}$ obtained by other experiments are also shown.
The limits on $A\rightarrow 2\gamma$ decay exclude a large new region of $g_{A\gamma}$ coupling constant ${(3\times 10^{-15}-10^{-8})~{}\rm{GeV}^{-1}}$ for the mass range $(0.02-5)$ MeV. The Borexino constrains are about 2-5 order of magnitude stronger than those obtained by laboratory-based experiments using nuclear reactors and accelerators and partly overlap the regions of heavy axion models Bere00 ; Bere01 ; Hal04 .
@@
$|g_{A\gamma}\times m_{A}^{3}|\leq 1.8\times 10^{-4}\rm{~{}eV^{2}~{}(\tan\beta=140)}$ (24)
@@
$|g_{A\gamma}\times m_{A}^{3}|\leq 2.2\times 10^{-3}\rm{~{}eV^{2}~{}(\tan\beta=0.25)}$ (26)
So, in the case of DFSZ-axion with $\tan\beta=140$ the coupling constant of 1 MeV axion is less than $g_{A\gamma}\leq 1.8\times 10^{-13}\rm{~{}GeV^{-1}}$.
@@
where $\sigma_{PC}$ is the Primakoff conversion cross sections; $N_{C}$ is the number of carbon nuclei in the 5 m radius sphere (M-C target), $T$ is live time of measurement and $\varepsilon_{PC}$ is the detection efficiency for 5.5 MeV $\gamma$’s (Fig.3 and table 1). The cross section $\sigma_{PC}$ is proportional to the $g^{2}_{A\gamma}$ and the axion flux $\Phi_{A}$ is proportional to the $g^{2}_{3AN}$ in accordance with (11) and (6). The number of Primakoff conversions $S_{PC}$ value depends on the product of the axion-photon and axion-nucleon coupling constants: $g^{2}_{A\gamma}\times g^{2}_{3AN}$. From the experimentally found relation $S_{PC}\leq S^{lim}_{PC}$ and for $(p_{A}/p_{\gamma})^{3}\approx 1$ one obtain the model-independent upper limit on the product of constants:
@@
$|g_{A\gamma}\times m_{A}|\leq 3.2\times 10^{-13}\rm{~{}(\tan\beta=140)}$ (29)
@@
$|g_{A\gamma}\times m_{A}|\leq 3.9\times 10^{-12}\rm{~{}(\tan\beta=0.25)}$ (31)
For DFSZ axion with $m_{A}$=1 MeV and $\tan\beta=140$, the upper limit on $g_{A\gamma}$ is correspond to $g_{A\gamma}\leq 3.2\times 10^{-10}\rm{~{}GeV^{-1}}$ that for a given mass is significantly weaker than the $A\rightarrow 2\gamma$ limit. But for axion masses $m_{A}$ less than $\simeq$50 keV, the limits obtained from the Primakoff conversion become more stringent. The regions of excluded values of $g_{A\gamma}$ and $m_{A}$ are shown in Fig.7 for three variants of axion models (three lines of contour 2).
The intersections of the KSVZ- and DFSZ-lines with the contours 1 and 2 in figure 7 defines the region of axion masses excluded by this experiment: for KSVZ-axion the region (0.7-80) keV and for DFSZ-axion $(\tan\beta=140)$ the region (1.2-100) keV are excluded, respectively. The constraints on the coupling constant $g_{A\gamma}$ are the most stringent in comparison with the other experiments in axion mass range (0.6-1.0) MeV; if the decay $A\rightarrow e^{+}e^{-}$ is not taken into account, the range is expanded to 5 MeV.
Figure 7: The limits on $g_{A\gamma}$ obtained by 1- present work ($A\rightarrow 2\gamma$, red line -DFSZ, $\tan\beta=140$, green - KSVZ, blue - DFSZ, $\tan\beta=0.25$), 2- present work (PC, the same notations as for 1), areas of excluded values are located inside contour, 3- CAST Zio05 ; Ari09 , 4 - HB Stars Raf08 , 5 - resonant absorption Gav18 , 6 - beam dump experiments Kon86 ; Bjo88 , 7- reactor experiment Alt95 ; Cha07 , 8 - expectation region from heavy axion models Bere00 ; Bere01 ; Hal04 . Some data taken from Oha23 .
@@
This constraint are completely model-independent and valid for any pseudoscalar particle. It’s important to stress that the limits were obtained on the assumption that axions escape from the Sun and reach the Earth, which implies $g_{Ae}<10^{-6}$ for $m_{A}<2m_{e}$ and $g_{Ae}<(10^{-11}-10^{-12})$ if $m_{A}>2m_{e}$ (Der10 ; Bel12a ).
Figure 8: The limits on the $g_{Ae}$ obtained by 1- present work (Compton process $A+e\rightarrow e+\gamma$, red line -DFSZ, $\tan\beta=140$, green - KSVZ, blue - DFSZ, $\tan\beta=0.25$), 2 - XENON1T Apr18 (relic axions), 3- Edelweiss Arm18 (relic axions), 4- read giants Raf08 , 5- PandaX-II Fu17 (relic and solar axions), 6- solar axion luminosity Gon09 , 7- ${\rm{{}^{83}Kr}}$ Gav22 (resonant absorption of solar axions) , 8- ${\rm{{}^{169}Tm}}$ Der09 ; Abd20 (resonant absorption of solar axions) , 9- reactor Alt95 ; Cha07 and solar experiments Bel08 ; Der10 , 10- beam dump experiments Kon86 ; Bjo88 , 11 - ortho-positronium decay Asa91 . Some data taken from Oha23 . The excluded values are located above the corresponding lines. The relations between $g_{Ae}$ and $m_{A}$ for DFSZ- and KSVZ-models are shown also.
@@
$|g_{Ae}\times m_{A}|\leq 2.7\times 10^{-6}~{}(\tan\beta=140)$ (34)
$|g_{Ae}\times m_{A}|\leq 4.8\times 10^{-6}~{}\rm{(KSVZ)}$ (35)
$|g_{Ae}\times m_{A}|\leq 3.3\times 10^{-5}~{}(\tan\beta=0.25)$ (36)
where $m_{A}$ is given in eV units. For DFSZ-axion with $m_{A}$=1 MeV and $\tan\beta=140$, the upper limit corresponds to $g_{Ae}\leq 2.7\times 10^{-12}$. The upper limit on KSVZ-axion mass is $m_{A}\leq 20$ keV and for DFSZ-axion $(\tan\beta=140)$ $m_{A}\leq 300$ eV.

### v2

@@ 4.2 Response function to axion reactions
The signature of these four reactions discussed above is a $5.5$ MeV peak. The Monte Carlo (MC) method based on the GEANT4 framework All16 was used to simulate the Borexino response to electrons and $\gamma$-quanta produced by axion interactions Ago18 . The simulations have taken into account the effect of ionization quenching, Cerenkov radiation and small non-linearity induced by the energy dependence on the event position.
@@
Table 1: The fraction $\epsilon$ of axion events surviving data selection procedure. $C$ - Compton axion to photon conversion, $A+e\rightarrow e+\gamma$; AE - axioelectric effect, $A+e+Z\rightarrow e+Z$; PC - Primakoff conversion on nuclei, $A+{{}^{12}\rm{C}}\rightarrow\gamma+{{}^{12}\rm{C}}$.
@@
Figure 3: Simulated responses to axion interactions in the Borexino IV: 1 - axioelectric effect ($5.49$ MeV electrons), 2 - Compton axion to photon conversion (electrons and $\gamma$-quanta), 3 - decay $A\rightarrow 2\gamma$, 4 - Primakoff conversion ($5.49$ MeV $\gamma$-quanta). The response functions are normalized to single axion interaction or decay in 55 m radius sphere.
@@ 4.3 Data selection
The Borexino detector carried out measurements from May 15, 2007 to October 7, 2021. The aim of data selection is to provide maximum exposure for the desired study with minimum background contribution. Background composition of the Borexino experiment was carefully studied in the course of many years of research. In the current analysis we aim to the energy region around of 5.5 MeV. The background components that have to be taken into account are the short-lived cosmogenic beta-sources ($^{12}\rm{B}$, 8He, 9C, 9Li, etc.), decays of internal 208Tl, external high-energy gamma produced due to neutron radiative capture in the SSS and solar neutrinos produced in the 8B decay.
@@
This system of cosmogenic veto allows for maximal suppression of the cosmogenic background for the price of 22.822.8% exposure loss.
Whilst the internal 208Tl and the solar neutrinos are uniform in the bulk of the detector and can not be suppressed or discriminated, the external high-energy gammas can be discriminated. Since these events count rate decreases exponentially towards the center of the detector, they can be suppressed by setting up a fiducial volume. In this analysis, the fiducial volume was defined by minimal distance to the inner vessel of $0.75$ m and had the mass of around $145$ tons. Moreover, these events were simulated with a full MC simulation following the detector evolution and this simulation was used for a so-called multivariate fit that takes into account the radial distribution providing a discrimination between external and internal spectrum components.
Figure 4: Energy spectra of the events and effect of the selection cuts. From top to bottom: 1 - raw spectrum; 2 - with 22 ms muon veto cut; 3 - final spectrum after application of the cosmogenic veto and FV cut. The bumps at $0.4$, $1.3$, $2.4$, and $3.3$ MeV are caused by decays of 210Po, 11C, external $\gamma$’s from 208Tl and internal 208Tl, respectively.
The experimental energy spectrum from Borexino in the range up to 6600 p.e. ($\simeq 12$ MeV), containing $579.3$ kton days of data, is shown in Fig. 4. At energies below $3$ MeV, the spectrum is dominated by $2.6$ MeV $\gamma$’s from the $\beta$-decay of 208Tl in the PMTs and in the SSS. The bumps in the spectrum observed at $200$ p.e., $700$ p.e., $1300$ p.e., and $1800$ p.e. are caused by $\alpha$-decays of 210Po, $\beta^{+}$-decays of $\rm{{}^{11}C}$, external $\gamma^{\prime}s$ from $\rm{{}^{208}Tl}$ and internal $\beta^{-}$-decays of $\rm{{}^{208}Tl}$, respectively (more details e.g. in Ago19 ).
@@ 5.1 Analysis of the spectrum
Figure 5 shows the observed Borexino energy spectrum in the (4.46.84.4-6.8) MeV ((22003400)(2200-3400) p.e.) range in which the axion peaks might appear. The spectrum $N(E)$ was fitted by the sum of four main components: the $\rm{{}^{208}Tl}$ decays inside the scintillator $N_{int}$, the simulated spectrum of external backgrounds caused by consequent $(\alpha,n)$ reactions followed by neutron captures $(n,\gamma)$ on Ni, Fe, Cr nuclei contained in the SSS constructions $N_{ext}$, the solar $\rm{{}^{8}B}$-neutrino-electron scattering events $N_{8B\nu}$, and the simulated spectrum of axion response $N_{A}$.
@@
Figure 5: Fit of the Borexino energy spectrum in the (4.46.84.4-6.8) MeV range. Curve 1 is the detector response function for axio-electric effect on carbon atoms and corresponds to 9090% c.l. upper limit ($S=5.6$ events). The fit (curve 2) is complementary to the radial fit on Fig. 6.
Figure 6: The fitted Borexino radial spectrum. The curve 1 corresponds to the uniform part of the spectrum whilst the curve 2 is the radial distribution of the simulated external background events. The decline in the distribution of uniform events is caused by the fiducial volume non-sphericity. The fit (curve 3) is complementary to the energy fit on Figure 5.
The limits obtained ($S^{\rm lim}_{avg}\simeq 1.7\times 10^{-3}$ counts/(100 t day) at 90% c.l.) are very low, e.g. $\sim 8\times 10^{4}$ times lower than expected number of events from solar pppp- neutrino (135 counts/(100 t day)). The upper limits on the number of events with energy 5.5 MeV constrain the product of axion flux $\Phi_{A}$ and the interaction cross section with electron, proton or carbon nucleus $\sigma_{A-e,p,C}$ via
@@
[…]
@@
5.2 Limits on $|g_{A\gamma}\times g_{3AN}|$-coupling
@@
The relationship between the isovector constant $g_{3AN}$ and the axion mass $m_{A}$ allows to constraint the $g_{A\gamma}$ depending on $m_{A}$. The three cases mentioned above were considered - the KSVZ-model with $E/N=0$ and the DFSZ-model for the angles $\tan\beta^{*}=140$ and $\tan\beta^{*}=0.25$ which are the boundary values of the allowed unknown angle $\beta^{*}$.
The excluded regions are inside three contours 1 in Fig. 7 (90 % c.l.). The curves from bottom to top correspond to DFSZ ($\tan\beta^{*}=140$), KSVZ and DFSZ ($\tan\beta^{*}=0.25$), respectively. For higher values of $g_{A\gamma}^{2}m_{A}^{3}$ axions decay before they reach the detector, while for lower $g_{A\gamma}^{2}m_{A}^{3}$ the probability of axion decay inside the Borexino volume is too low. The limits on $g_{A\gamma}$ obtained by other experiments are also shown.
The limits on $A\rightarrow 2\gamma$ decay exclude a large new region of $g_{A\gamma}$ coupling constant $(3\times 10^{-15}-10^{-8})~\rm{GeV}^{-1}$} for the mass range (0.025)(0.02-5) MeV. The Borexino constrains are about 2-5 orders of magnitude stronger than those obtained by laboratory-based experiments using nuclear reactors and accelerators and partly overlap the regions of heavy axion models Bere00 ; Bere01 ; Hal04 .
@@
$|g_{A\gamma}\times m_{A}^{3}|\leq 1.8\times 10^{-4}\rm{~eV^{2}~(\tan\beta^{*}=140)}$ (24)
@@
$|g_{A\gamma}\times m_{A}^{3}|\leq 2.2\times 10^{-3}\rm{~eV^{2}~(\tan\beta^{*}=0.25)}$ (26)
So, in the case of DFSZ-axion with $\tan\beta^{*}=140$ the coupling constant of 1 MeV axion is less than $g_{A\gamma}\leq 1.8\times 10^{-13}\rm{~GeV^{-1}}$.
@@
where $\sigma_{PC}$ is the Primakoff conversion cross sections; $N_{C}$ is the number of carbon nuclei in the 5 m radius sphere (Monte-Carlo (M-C) target), $T$ is live time of measurement and $\varepsilon_{PC}$ is the detection efficiency for 5.5 MeV $\gamma$’s (Fig.3 and table 1). The cross section $\sigma_{PC}$ is proportional to the $g^{2}_{A\gamma}$ and the axion flux $\Phi_{A}$ is proportional to the $g^{2}_{3AN}$ in accordance with (11) and (6). The number of Primakoff conversions $S_{PC}$ value depends on the product of the axion-photon and axion-nucleon coupling constants: $g^{2}_{A\gamma}\times g^{2}_{3AN}$. From the experimentally found relation $S_{PC}\leq S^{lim}_{PC}$ and for $(p_{A}/p_{\gamma})^{3}\approx 1$ one obtain the model-independent upper limit on the product of constants:
@@
$|g_{A\gamma}\times m_{A}|\leq 3.2\times 10^{-13}\rm{~(\tan\beta^{*}=140)}$ (29)
@@
$|g_{A\gamma}\times m_{A}|\leq 3.9\times 10^{-12}\rm{~(\tan\beta^{*}=0.25)}$ (31)
For DFSZ axion with $m_{A}$=1 MeV and $\tan\beta^{*}=140$, the upper limit on $g_{A\gamma}$ corresponds to $g_{A\gamma}\leq 3.2\times 10^{-10}\rm{~GeV^{-1}}$ that for $m_{A}$=1 MeV is significantly weaker than the $A\rightarrow 2\gamma$ limit. But for axion masses $m_{A}$ less than $\simeq$50 keV, the limits obtained from the Primakoff conversion become more stringent. The regions of excluded values of $g_{A\gamma}$ and $m_{A}$ are shown in Fig.7 for three variants of axion models (three lines of contour 2).
The intersections of the KSVZ- and DFSZ-lines with the contours 1 and 2 in figure 7 defines the region of axion masses excluded by this experiment: for KSVZ-axion the region (0.7-80) keV and for DFSZ-axion $(\tan\beta^{*}=140)$ the region (1.2-100) keV are excluded, respectively. The constraints on the coupling constant $g_{A\gamma}$ are the most stringent in comparison with the other experiments in axion mass range (0.6-1.0) MeV; if the decay $A\rightarrow e^{+}e^{-}$ is not taken into account, the range is expanded to 5 MeV.
Figure 7: The limits on $g_{A\gamma}$ obtained by 1 - present work ($A\rightarrow 2\gamma$-decay (from bottom to top: DFSZ ($\tan\beta^{*}=140$, red line), KSVZ (green), DFSZ ($\tan\beta^{*}=0.25$, blue))), 2 - present work (PC, the same notations as for 1), areas of excluded values are located inside contour, 3- CAST Zio05 ; Ari09 , 4 - HB Stars Raf08 , 5 - resonant absorption Gav18 , 6 - beam dump experiments Kon86 ; Bjo88 , 7- reactor experiments Alt95 ; Cha07 , 8 - expectation region from heavy axion models Bere00 ; Bere01 ; Hal04 . Some data taken from Oha23 . The relations between $g_{Ae}$ and $m_{A}$ for DFSZ- and KSVZ-models are shown also.
@@
This constraint is completely model-independent and valid for any pseudoscalar particle. The previous limit is improved by a factor of $\sim 3$ Bel12a , but one must take into account the complicating fact that the expected number of axion events is proportional to the square of the product $g_{Ae}$ and $g_{3AN}$ constants: $S_{CC}\sim|g_{Ae}\times g_{3AN}|^{2}$. It’s important to stress that the limits were obtained on the assumption that axions escape from the Sun and reach the Earth, which implies $g_{Ae}<10^{-6}$ for $m_{A}<2m_{e}$ and $g_{Ae}<(10^{-11}-10^{-12})$ if $m_{A}>2m_{e}$ (Der10 ; Bel12a ).
Figure 8: The limits on the $g_{Ae}$ obtained by 1 - present work (Compton process $A+e\rightarrow e+\gamma$ (from bottom to top: DFSZ ($\tan\beta^{*}=140$, red line), KSVZ (green), DFSZ ($\tan\beta^{*}=0.25$, blue)), 2 - XENON1T Apr18 (relic axions), 3 - Edelweiss Arm18 (relic axions), 4 - red giants Raf08 , 5- PandaX-II Fu17 (relic and solar axions), 6 - solar axion luminosity Gon09 , 7 - $\rm{{}^{83}Kr}$} Gav22 (resonant absorption of solar axions) , 8 - $\rm{{}^{169}Tm}$} Der09 ; Abd20 (resonant absorption of solar axions) , 9 - reactor Alt95 ; Cha07 and solar experiments Bel08 ; Der10 , 10 - beam dump experiments Kon86 ; Bjo88 , 11 - ortho-positronium decay Asa91 . Some data taken from Oha23 . The relations between $g_{Ae}$ and $m_{A}$ for DFSZ- and KSVZ-models are shown also.
@@
$|g_{Ae}\times m_{A}|\leq 2.7\times 10^{-6}~\rm{eV}~(\tan\beta^{*}=140)$ (34)
$|g_{Ae}\times m_{A}|\leq 4.8\times 10^{-6}~\rm{eV}~\rm{(KSVZ)}$ (35)
$|g_{Ae}\times m_{A}|\leq 3.3\times 10^{-5}~\rm{eV}~(\tan\beta^{*}=0.25)$ (36)
where $m_{A}$ is given in eV units. For DFSZ-axion with $m_{A}$=1 MeV and $\tan\beta^{*}=140$, the upper limit corresponds to $g_{Ae}\leq 2.7\times 10^{-12}$. The upper limit on KSVZ-axion mass is $m_{A}\leq 20$ keV and for DFSZ-axion $(\tan\beta^{*}=140)$ - $m_{A}\leq 300$ eV.